OpenStudy (anonymous):

How do I find the unit rate in feet per second for 300 yards/ 1 minute?

OpenStudy (mathmale):

In this case, rate =distance/time. We have the rate 300 yds/1 min. First, convert yards to feet. Second, convert minutes to seconds.

OpenStudy (anonymous):

How many feet are in a yard ?

OpenStudy (anonymous):

Okay so I know that I have 900 feet, since 3*300 is 900. But I don't know how to convert the minutes to seconds. I know there are 60 seconds in a minute.

OpenStudy (anonymous):

thats it....900 ft/ 60 sec

OpenStudy (anonymous):

 that's what I have written down on my paper so far.

OpenStudy (anonymous):

so it IS 900 ft/60 sec?

OpenStudy (anonymous):

I am not sure...because it is asking feet per second...not per 60 seconds

OpenStudy (anonymous):

900/60 = x/1 (900 ft to 60 sec = x ft to 1 sec) cross multiply 60x = 900 x = 900/60 x = 15 15 ft per second

OpenStudy (anonymous):

unit rate is 15 ft per second

OpenStudy (anonymous):

you had it you just didn't finish it